OpenStudy (nannie):

Here is another. Please help! Divide and Simplify 11z-66/4÷z-6/6z

OpenStudy (anonymous):

\[\frac{44z-66}{4}*\frac{6z}{6z^2-6}\] \[\frac{22(2z-3)}{2}*\frac{3z}{6(z^2-1)}\] \[11(2z-3)*\frac{1z}{2(z^2-1)}\] \[\frac{11z(2z-3)}{2(z^2-1)}\] \[\frac{22z^2-33z}{2z^2-2}\]